OpenStudy (anonymous):

The function f is defined by f(x)= sqrt (25-x^2) for -5 less than/equal to x less than/equal to 5 a. write an equation for the line tangent to the graph of f at x=-3

OpenStudy (anonymous):

take the derivative, replace \(x\) bu \(-3\)

OpenStudy (anonymous):

\[f(x)=\sqrt{25-x^2}\] \[f'(x)=\frac{-x}{\sqrt{25-x^2}}\]

OpenStudy (anonymous):

i am going to guess you get \(\frac{3}{4}\) but you can check it

OpenStudy (anonymous):

what i mean is that the slope will be \(\frac{3}{4}\)

OpenStudy (anonymous):

it is clear that this function is the upper half of the circle with radius 5 centered at the origin?

OpenStudy (anonymous):

It's not specified in prompt if that's what you mean, but i guess it would be